OpenStudy (anonymous):

∫ cos x / (sin² x + 2sinx) dx you must use tables in back of calculus book.

hartnn (hartnn):

plug in u = sin x

hartnn (hartnn):

then du = ... ?

hartnn (hartnn):

@Josh_S , you there buddy ?

OpenStudy (anonymous):

im here

OpenStudy (anonymous):

i understand how to do it that way, but this section requires you to use the tables in the back of the book. im not sure which table to use.

OpenStudy (anonymous):

sorry if you dont have a book handy

hartnn (hartnn):

then i won't know which formulas r there in the back...and hence which to use.....do u have any guesses on which can be used ?

OpenStudy (anonymous):

i could take pictures of the tables and attach the files?

hartnn (hartnn):

yes, you could. but that would take time, if you have, go on

OpenStudy (anonymous):

i tried uploading pictures it didnt work

hartnn (hartnn):

that.is.a.long.list -_- searching....

OpenStudy (anonymous):

sorry, thanks!

hartnn (hartnn):

gotcha! 49 !

hartnn (hartnn):

u (a+bu)

OpenStudy (anonymous):

i wanted one that was du/u^2 + 2u but theres not any

hartnn (hartnn):

after plugging in sin x = u we get, in the denominator as u^2+2u = u (2+u) a=2, b=1 got this ?

OpenStudy (anonymous):

what number is that one im looking for it

hartnn (hartnn):

look for 49 th !

OpenStudy (anonymous):

you sir are a genuis

hartnn (hartnn):

lol, no sir! i just have good amount of practice :) you practice more, you'll be a greater genius than me ;)

OpenStudy (anonymous):

thanks!!! im about to post another question in a minute if you want to tackle that one too. :)

hartnn (hartnn):

i would like to :) i love integration!